OpenStudy (anonymous):

Write the equation of the quadratic function with roots -10 and -8 and a vertex at (-9, -3). Write the equation of the quadratic function with roots 0 and 2 and a vertex at (1, 5). f Write the equation of the quadratic function with roots 6 and 10 and a vertex at (8, 2 Somebody show me the steps for these please. Its been explained already BUT I DONT UNDERSTAND IT. Please :/

OpenStudy (anonymous):

For "Write the equation of the quadratic function with roots 6 and 10 and a vertex at (8, 2)" a(x-6)(x-10) = a(x-8)^2 +2 Solve this for a.

OpenStudy (anonymous):

in order to get the point that fct corss the xy axis, we have to know that -b/2a=-9, and let f=ax^2+bx+c. then x1=-b+derta^0.5/2a x1=-b-derta^0.5/2a and derta=b^2-4ac then u can get all of them

OpenStudy (anonymous):

like the first one is fx=x^2+18x+84

OpenStudy (anonymous):

sry! its fx=x^2+18x+78

OpenStudy (anonymous):

Thank you.